OpenStudy (anonymous):

A limit (challenge) problem for smart people. In the question n is an integer Condition:- Try to solve it in not more than 3-4 steps.

ganeshie8 (ganeshie8):

Is \(n\) an integer ?

OpenStudy (anonymous):

Yes n is an integer. :)

OpenStudy (dan815):

0

OpenStudy (anonymous):

no

OpenStudy (dan815):

infinity

OpenStudy (dan815):

1

OpenStudy (anonymous):

no

OpenStudy (dan815):

pi

OpenStudy (dan815):

e,

OpenStudy (anonymous):

u need to show your method , don't just give out answers like a idiot

OpenStudy (dan815):

lulz :)

OpenStudy (dan815):

maybe complex numbers i dont like the root in there

OpenStudy (anonymous):

no real sollution :)

OpenStudy (dan815):

you can get real solution from that

OpenStudy (dan815):

is it indeterminate?

OpenStudy (dan815):

is the answer even converging

OpenStudy (anonymous):

no there is a answer

OpenStudy (dan815):

okay then ill work on it : )

OpenStudy (anonymous):

It dosen't have stupid answers like it does not exist etc

OpenStudy (trojanpoem):

No.name is right. It got answer.

OpenStudy (anonymous):

You got the answer @TrojanPoem

OpenStudy (anonymous):

?

OpenStudy (trojanpoem):

Nearly , but more than 3 steps.

OpenStudy (anonymous):

its ok tell your answer

ganeshie8 (ganeshie8):

im getting pi but im still messing wid it

OpenStudy (anonymous):

@ganeshie8 you are correct !

OpenStudy (dan815):

psh i already guessed pi lol

OpenStudy (anonymous):

Ok , before i post my sollution , how did u get it @ganeshie8

ganeshie8 (ganeshie8):

bit embarrassing i kno 

OpenStudy (trojanpoem):

haha @ganeshie8 Like ++

OpenStudy (anonymous):

You made my day @ganeshie8 :)

ganeshie8 (ganeshie8):

I'm glad you find it was entertaining :) please provide a hint as im kinda stuck..

OpenStudy (anonymous):

no problem , never mind Well the hint is itself the whole sollution , so let me post the sollution itself ok?

OpenStudy (dan815):

no odont

ganeshie8 (ganeshie8):

^ give us some time

OpenStudy (anonymous):

Yeah don't give up so easily , u will get it!

OpenStudy (anonymous):

After giving hint it is easy , so i will not give any hint :)

OpenStudy (anonymous):

Ok a small hint

ganeshie8 (ganeshie8):

nope wait almost there

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

Don't even think of binomial expansions , you will get the answer no doubt , but that's a waste of time and space

ganeshie8 (ganeshie8):

\[\large{\begin{align}&\lim\limits_{n\to\infty}~n*\sin(2\pi\sqrt{n^2+1})\\~\\ &=\lim\limits_{n\to\infty}~n*\sin(2\pi(\sqrt{n^2+1}-n)+2n\pi)\\~\\ &=\lim\limits_{n\to\infty}~n*\sin(2\pi(\sqrt{n^2+1}-n))\\~\\ &=\lim\limits_{n\to\infty}~n*\sin\left(\frac{2\pi}{\sqrt{n^2+1}+n}\right)\\~\\ \end{align}}\]

ganeshie8 (ganeshie8):

next i want to use sinx/x limit but im gonna need more paper, one sec..

OpenStudy (dan815):

its so weird i cant believe the 1 is skewing it form 2pi to pi

OpenStudy (dan815):

for a big number sqrt(1+n^2) is pretty much sqrt(n^2)

OpenStudy (dan815):

so i said it was just n*sin(2pi*n) and thats inf * 0 so from lohopitals ud get 2pi

OpenStudy (dan815):

but that +1 in there changed it to pi O_O still working

ganeshie8 (ganeshie8):

I just figured out a really neat way to solve this w/o using lhopital xD

OpenStudy (dan815):

are you squaring and sqrting the whole limit

OpenStudy (dan815):

or k^2=1+n^2!

ganeshie8 (ganeshie8):

nvm its not working, was trying to interpret the given expression as a variation of area of regular polygon formula... but no success

ganeshie8 (ganeshie8):

as \(n\to\infty\), the regular polygon becomes an unit circle evaluating the limit to \(\pi\) thought it would be that simple but looks tricky to conclude like that

OpenStudy (dan815):

ya not sure what im doing wrong, i keep getting 2pi again

OpenStudy (dan815):

look at my work, can u spot something wrong?

ganeshie8 (ganeshie8):

okie

OpenStudy (dan815):